OpenStudy (anonymous):

simplify: sin[(3pi over 2)+x)]

OpenStudy (anonymous):

Use the sum of angles identity for sine: sin(a+b) = (sin a)(cos b) + (cos a)(sin b) sin(3π/2+x) = (sin 3π/2)(cos x) + (cos 3π/2)(sin x) = (-1)(cos x) + (0)(sin x) = -cos x

OpenStudy (dumbcow):

what they said ^^

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

@camjan Welcome:)

OpenStudy (anonymous):

hehe.. can i add u on facebook?

OpenStudy (anonymous):

@camjan yeah, sure:)